Question 1. Let u = (2,1,0,5) e R2. Apply the process of inner product spaces.

  • Solve u for its length
  • Solve u for its Normalized vector
1
Expert's answer
2022-02-01T16:37:24-0500

1] The length of u is;

"\displaystyle
\|u\|=\sqrt{u\cdot u}=[(2,1,0,5)\cdot(2,1,0,5)]^{\frac{1}{2}}=[2\times2+1\times1+0\times0+5\times5]^{\frac{1}{2}}\\
\qquad\ =[4+1+0+25]^{\frac{1}{2}}=\sqrt{30}"


2] The normalized vector of u is;

"\displaystyle
\frac{u}{\|u\|}=\frac{1}{\sqrt{30}}(2,1,0,5)"